OpenStudy (anonymous):

Solve for x. -3x - 8 + 6x = 7

OpenStudy (saifoo.khan):

Add like terms, (-3+6)x = 7 + 8 3x = 15 x = 15/3 x = 5

OpenStudy (anonymous):

-_-

OpenStudy (saifoo.khan):

Slow people these days. Good for nothing.

OpenStudy (anonymous):

Hmm..fast people these days. Too quick for the world.

OpenStudy (saifoo.khan):

The faster the better.

OpenStudy (anonymous):

Slow and steady wins the race.

OpenStudy (saifoo.khan):

LOoooooool. nice joke.

OpenStudy (anonymous):

No joke. Go back to kindergarten.

OpenStudy (saifoo.khan):

Look who's speaking?

OpenStudy (anonymous):

Whos speaking? I dont hear anyone O.o

OpenStudy (saifoo.khan):

better get your ears cleaned.

OpenStudy (anonymous):

No, they clean pretty well.

OpenStudy (saifoo.khan):

o_O

OpenStudy (anonymous):

Yes. They are. I clean my ears. You must wear your glasses.

OpenStudy (saifoo.khan):

im wearing mine already. thankyou.

OpenStudy (anonymous):

Well, get a thicker size. You really need it.

OpenStudy (saifoo.khan):

you aren't an online doctor.

OpenStudy (anonymous):

Who knows...i just may be one.

OpenStudy (saifoo.khan):

i know you aren't.

OpenStudy (anonymous):

Where is your proof?

OpenStudy (anonymous):

You cant just "know things"

OpenStudy (saifoo.khan):

how are you so sure i can't know things? where's your proof?

OpenStudy (saifoo.khan):

Some things are kept secretive and the proves aren't revealed. and it's good.

OpenStudy (anonymous):

Because your just some guy that comes here. And i dont even know you, nor do you know me. Theres my proof. Wheres yours?

OpenStudy (saifoo.khan):

That isn't a proof. and i know you. idk if you forgot me.. you need to see a doc. you might be suffering from short-term memory less.

OpenStudy (anonymous):

Nooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooo. You lie. Dont say nothing else. But even so, it doesnt matter, you STILL dont knw me in person, or seen how i look like. So what the heck.

OpenStudy (saifoo.khan):

Today, world is virtual. knowing someone in real life dosn't really matter.

OpenStudy (saifoo.khan):

"until it's next level"

OpenStudy (anonymous):

Hmm..but does the virtual world really what people think it is?

OpenStudy (saifoo.khan):

is that an incomplete statement?

OpenStudy (anonymous):

No. Its pretty complete.

OpenStudy (anonymous):

I stay true to myself. But do others?

OpenStudy (saifoo.khan):

im getting a feeling it's incomplete.

OpenStudy (saifoo.khan):

Sure, others are truthful too.

OpenStudy (anonymous):

But how can a person really know?

OpenStudy (saifoo.khan):

know about..?

OpenStudy (anonymous):

If a person is really true to them selves in the virtual world?

OpenStudy (saifoo.khan):

Ah. People get the feeling already by the way they talk.

OpenStudy (anonymous):

I guess you are right.

OpenStudy (saifoo.khan):

Like always. ^_~

OpenStudy (anonymous):

NOT Always. More like, SOMETIMES.

OpenStudy (saifoo.khan):

It's always!

OpenStudy (anonymous):

NODODY IS PERFECT.

OpenStudy (saifoo.khan):

im Nobody. ;)

OpenStudy (anonymous):

Then who the hell are you?

OpenStudy (saifoo.khan):

im Nobody, so im perfect.

OpenStudy (anonymous):

Hate to break your bubble. But if you are "Nobody" Then you cant be "ANYBODY" which means you dont even exist, which means your not perfect.

OpenStudy (saifoo.khan):

im not anybody. im nobody. there's a difference between both.

OpenStudy (anonymous):

Show difference.

OpenStudy (saifoo.khan):